# Closed-Form Feedback-Free Learning with Forward Projection

### v2

@@ A.9 Optimisation Theory
We now compare the fit of the Forward Projection estimator in a two-layer network, to that of a simple linear estimator in random data. We consider a dataset $\{\textbf{x}_{i},\textbf{y}_{i}\}_{i=1}^{N}$ with random Gaussian distribution such that $\textbf{x}_{i}\in\mathcal{N}(\textbf{0},\textbf{I}_{m})$ and an independent random normal target $\textbf{y}_{i}\in\mathcal{N}(\textbf{0},\textbf{I}_{d})$. We assume $N>m>d$. We consider a fully connected neural network with a single hidden layer of dimension $m$, with relu-activation. The task is to identify network weights $\{\textbf{W}_{1},\textbf{W}_{2}\}$ to minimise the error $\lVert\textbf{Y}-\operatorname{ReLU}(\textbf{XW}_{1})\textbf{W}_{2}\rVert$. We first consider the expected error of a single layer, linear model. For convenience, we will set the penalty factor $\lambda=0$, such that
@@
The idempotent matrix $(\textbf{I}-\textbf{P}_{\textbf{x}})$ has rank NmN-m, as it a projection matrix onto the orthogonal complement of the column space of X. Therefore, the expected squared Frobenius norm of the error is given by
@@
$\mathbb{E}[\textbf{X}\textbf{W}_{1}\textbf{U}^{+}]=\mathbb{E}[\textbf{XQU}^{+}]+\mathbb{E}[\textbf{P}_{\textbf{x}}\textbf{YUU}^{+}]=\textbf{P}_{\textbf{x}}\textbf{Y}=\frac{1}{2}\hat{\textbf{Y}}_{1}.$ (38)
@@
As $m>d$ and U has random Gaussian entries, we have $\operatorname{Var}[\operatorname{tr}(\textbf{U}^{T})]=d$. Therefore, we have
@@
Therefore, even when modelling pure Gaussian noise, the error of the two-layer model is lower than that of the simple linear predictor.
A.10 Modelling high dimensional labels
@@
Therefore we have $\mathbb{E}[\operatorname{Tr}\left((\textbf{U}_{l}\textbf{U}_{l}^{T})^{-1}\right)]=\frac{m_{L}}{m_{l}-m_{L}-1}$ for $m_{L}<m_{l}$. Thus, the expectation $\mathbb{E}\left[\operatorname{Tr}\left(\textbf{U}_{l}^{+}\left(\textbf{U}_{l}^{+}\right)^{T}\right)\right]$ increases with $m_{L}$ and decreases with $m_{l}$. Thus, higher dimensional labels are subject to poorer encoding, for a given hidden layer dimension. Conversely, for a given label dimension, it is expected that representation and learning will improve with increases of the hidden layer dimension. Accordingly, adequate modelling of high dimensional labels may require larger hidden layer dimensions.
A.11 Modelling with Challenging Hidden Activation Functions

### v3

@@ A.9 Optimisation Theory
We now compare the fit of the Forward Projection estimator in a two-layer network, to that of a simple linear estimator in random data. Consider a dataset $\{\textbf{x}_{i},\textbf{y}_{i}\}_{i=1}^{N}$ drawn from a multivariate Gaussian distribution, where $\textbf{x}_{i}\sim\mathcal{N}(\textbf{0},\textbf{I}_{m})$ and the independent target $\textbf{y}_{i}\sim\mathcal{N}(\textbf{0},\textbf{I}_{d})$. Assume $N>m>d$. We consider a fully connected neural network with a single hidden layer of dimension $m$, with ReLU activation. The task is to identify network weights $\{\textbf{W}_{1},\textbf{W}_{2}\}$ to minimise the error $\lVert\textbf{Y}-\operatorname{ReLU}(\textbf{XW}_{1})\textbf{W}_{2}\rVert$. We first consider the expected error of a single layer, linear model. For convenience, we will set the penalty factor $\lambda=0$, such that
@@
The idempotent matrix $(\textbf{I}-\textbf{P}_{\textbf{x}})$ has rank NmN-m, as it is a projection onto the orthogonal complement of the column space of X. Therefore, the expected squared Frobenius norm of the error is given by
@@
$\mathbb{E}[\textbf{X}\textbf{W}_{1}\textbf{U}^{+}]=\mathbb{E}[\textbf{XQU}^{+}]+\mathbb{E}[\textbf{P}_{\textbf{x}}\textbf{YUU}^{+}]=\textbf{P}_{\textbf{x}}\textbf{Y}=\hat{\textbf{Y}}_{1}.$ (38)
@@
As $m>d$ and U has random standard Gaussian entries, we have $\operatorname{Var}[\operatorname{tr}(\textbf{U}^{T})]=d$. Therefore, we have
@@
Thus, even under pure Gaussian noise, the two-layer model achieves a lower expected error than the simple linear predictor.
A.10 Iterative learning with Forward Projection
An alternative to deriving the closed-form Forward Projection solution for regression over the entire dataset is to approximate the optimization procedure through iterative updates on mini-batches, while preserving local learning dynamics and target generation as formalized in Equation 4. We partition the dataset $\mathcal{D}$ into $T$ mini-batches $\mathcal{B}_{t}=\{(\mathbf{x}_{j},\mathbf{y}_{j})\}_{j=1}^{B}$, where $b$ is the batch size. For each mini-batch, compute $\mathbf{A}_{l-1}^{(t)}\in\mathbb{R}^{B\times m_{l-1}}$ and $\tilde{\mathbf{Z}}_{l}^{(t)}\in\mathbb{R}^{B\times m_{l}}$ using the target generation rule
$\tilde{\mathbf{z}}_{l}=g_{l}(\mathbf{a}_{l-1}\mathbf{Q}_{l})+g_{l}(\mathbf{y}\mathbf{U}_{l}).$ (46)
For layer $l$, define the local loss:
$\mathcal{L}_{l}^{(t)}=\frac{1}{B}\sum_{j=1}^{B}\|\mathbf{z}_{j,l}-\tilde{\mathbf{z}}_{j,l}\|^{2},$ (47)
where $\mathbf{z}_{l}^{(j)}=\mathbf{a}_{l-1}^{(j)}\mathbf{W}_{l}$. The gradient of $\mathcal{L}_{l}^{(t)}$ with respect to $\mathbf{W}_{l}$ is:
$\nabla_{\mathbf{W}_{l}}\mathcal{L}_{l}^{(t)}=\frac{2}{B}(\mathbf{A}_{l-1}^{(t)})^{\top}(\mathbf{A}_{l-1}^{(t)}\mathbf{W}_{l}-\tilde{\mathbf{Z}}_{l}^{(t)}).$ (48)
Therefore, weights may be updated according to
$\mathbf{W}_{l}\leftarrow\mathbf{W}_{l}-\eta\nabla_{\mathbf{W}_{l}}\mathcal{L}_{l}^{(t)},$ (49)
where $\eta$ is the learning rate. Optionally, regularisation may be included via the ridge penalty:
$\nabla_{\mathbf{W}_{l}}\mathcal{L}_{l}^{(t)}\;+=\;2\lambda\mathbf{W}_{l}.$ (50)
In experiments on FMNIST classification by MLPs with $3\times 1000$ hidden neurons, sequential FP training achieved similar test performance to the closed-form FP solution (test AUC: $98.6\pm 0.79\%$, test accuracy: $84.4\pm 0.02\%$).
A.11 Modelling high dimensional labels
@@
Therefore we have $\mathbb{E}[\operatorname{Tr}\left((\textbf{U}_{l}\textbf{U}_{l}^{T})^{-1}\right)]=\frac{m_{L}}{m_{l}-m_{L}-1}$ for $m_{L}+1<m_{l}$. Thus, the expectation $\mathbb{E}\left[\operatorname{Tr}\left(\textbf{U}_{l}^{+}\left(\textbf{U}_{l}^{+}\right)^{T}\right)\right]$ increases with $m_{L}$ and decreases with $m_{l}$. Thus, higher dimensional labels are subject to poorer encoding, for a given hidden layer dimension. Conversely, for a given label dimension, it is expected that representation and learning will improve with increases of the hidden layer dimension. Accordingly, adequate modelling of high dimensional labels may require larger hidden layer dimensions.
A.12 Training complexity of other methods
In this section, we describe the training complexity involved in each step of the different methods that are used in this paper for benhcmarking purposes.
A.12.1 Backpropagation
Backpropagation requires a forward pass through each layer ($N_{e}Nm^{2}$), a backward pass to compute weight gradient ($N_{e}Nm^{2}$) and the input gradient ($N_{e}Nm^{2}$). After each batch, weights are updated ($N_{e}B^{-1}Nm^{2}$). For each layer, backpropation requires storage of weight parameters ($m^{2}$), input activations ($Bm$), and weight gradients ($m^{2}$).
A.12.2 Local Supervision
In local supervision, after the forward pass through the main layer ($N_{e}Nm^{2}$ MACs), an auxiliary forward pass is conducted through the local supervision head ($N_{e}Nmm_{L}$ MACs). Subsequently, the auxiliary gradient is computed at the local supervision head ($N_{e}Nmm_{L}$ MACs), which is backpropagated through the main layer ($N_{e}Nm^{2}$ MACs). Forward weight updates for the main layer occur at the end of each batch ($N_{e}B^{-1}Nm^{2}$ MACs). Optionally, the auxiliary layer may be trained ($N_{e}B^{-1}Nmm_{L}$ MACs). Parameters are stored for the main layer weights ($m^{2}$) and auxiliary weights $mm_{L}$. Local supervision requires storage of main layer weights ($m^{2}$), auxiliary layer weights ($mm_{L}$), main layer input activations ($Bm$), main layer output activations ($Bm$), auxiliary layer output activations ($Bm_{L}$) and weight updates ($m^{2}$).
A.12.3 Forward-Forward (FF)
For a given layer, FF requires two forward passes – for “positive” and “negative” data, respectively ($2N_{e}Nm^{2}$ MACs) [15]. A goodness metric is computed between positive and negative activations, requiring $N_{e}Nm$ MACs. Local gradients are computed for both positive and negative passes ($2N_{e}Nm^{2}$ MACs). Forward weight updates occur with each batch ($B^{-1}N_{e}Nm^{2}$ MACs). At inference time, forward passes are required for each possible class label ($Nm^{2}m_{l}$ MACs). Forward-Forward requires storage of weight parameters ($m^{2}$), positive and negative input activations ($2B(m+m_{L})$), positive and negative output activations ($2Bm$) and weight updates ($m^{2}$).
A.12.4 Difference Target Propagation (DTP)
DTP replaces gradient transport with targets formed using a learned approximate inverse (decoder) and a difference correction. Training requires an encoder forward pass ($N_{e}Nm^{2}$ MACs), and two decoder passes ($2N_{e}Nm^{2}$ MACs) to perform the difference correction operation between targets and observed activations [1]. Decoder training employs a denoising style loss, forwarding through both encoder ($N_{e}Nm^{2}$ MACs) and decoder ($N_{e}Nm^{2}$ MACs), before computing the decoder weight gradient ($N_{e}Nm^{2}$ MACs). Finally, the encoder weight gradient is computed with respect to the local target loss ($N_{e}Nm^{2}$ MACs). Weight updates for both encoder and decoder occur after each batch ($2B^{-1}N_{e}Nm^{2}$ MACs). At inference time, only the encoder forward pass is required. Assuming batch size $B=1$, training requires a forward pass ($N_{e}Nm^{2}$ MACs) and $8N_{e}Nm^{2}$ further MACs per layer. DTP requires storage of encoder and decoder weight parameters ($2m^{2}$), input activations ($Bm$), output activations ($Bm$), targets ($Bm$) and weight updates ($m^{2}$), which can be performed sequentially for encoder and decoder.
A.12.5 Predictive Coding
PC performs $k$ iterations of activity inference that alternate forward prediction ($kN_{e}Nm^{2}$ MACs) and feedback error projection through a decoder network ($kN_{e}Nm^{2}$ MACs), followed by a local Hebbian weight gradient computation ($N_{e}Nm^{2}$ MACs) [5]. Forward weight updates occur at the end of each batch ($N_{e}B^{-1}Nm^{2}$ MACs). Letting $k=1$, $B=1$, training requires a forward pass ($N_{e}Nm^{2}$ MACs) and $3N_{e}N_{m}^{2}$ further MACs per layer. Predictive coding requires storage of weight parameters ($m^{2}$), input activations ($m$), output activations ($Bm$), output targets ($Bm$) and hebbian weight updates ($m^{2}$).
A.13 Modelling with Challenging Hidden Activation Functions